# Non-Hermitian symmetry breaking and Lee-Yang theory for quantum XYZ and clock models

### v1

@@ IV Conclusion
In summary, we have generalized the Lee-Yang theory based on fidelity zeros to both the XYZ model and the $\mathbb{Z}_{3}$ clock model. The fidelity zeros in these systems, arising from non-Hermitian symmetry breaking, exhibit distinct distributions in the ordered and disordered phases, consistently obey the Lee-Yang theorem, and faithfully signal the phase transition. Our results show that this approach is not only effective for the anisotropic XYZ model but also capable of capturing criticality in systems with higher discrete symmetries, such as the $\mathbb{Z}_{3}$ clock model. Consequently, the fidelity-based Lee-Yang framework provides a simpler implementation with robust numerical accuracy for identifying quantum phase transitions, thereby opening avenues for applying the Lee-Yang formalism to a broader class of correlated quantum systems. A promising direction for future work is to explore its extension to continuous symmetry-breaking cases, such as U(1) and SU(2).

### v2

@@ IV Conclusion
In summary, we have generalized the Lee-Yang theory based on fidelity zeros to both the XYZ model and the $\mathbb{Z}_{3}$ clock model. The fidelity zeros in these systems, arising from non-Hermitian symmetry breaking, exhibit distinct distributions in the ordered and disordered phases, consistently obey the Lee-Yang theorem, and faithfully signal the phase transition. Our results show that this approach is not only effective for the anisotropic XYZ model but also capable of capturing criticality in systems with higher discrete symmetries, such as the $\mathbb{Z}_{p}$ clock model. Consequently, the fidelity-based Lee-Yang framework provides a simpler implementation with robust numerical accuracy for identifying quantum phase transitions, thereby opening avenues for applying the Lee-Yang formalism to a broader class of correlated quantum systems. A promising direction for future work is to explore its extension to continuous symmetry-breaking cases, such as U(1) and SU(2) symmetries.